# Preconditioning without a preconditioner using randomized block Krylov subspace methods

### v2

@@ 6.1 Sampling with the inverse covariance
Computing $\mathbf{A}^{-1/2}\mathbf{b}$ is used to move to a “whitened” coordinate space, and has found use in a number of data-science applications [37, 52].
We can define an approximation for applying inverse square roots similar to the block-Lanczos square root iterate (Definition 6.2).
@@
Note that $\mathsf{bisq}_{t}^{(i)}=\mathbf{A}^{-1}\mathsf{bsq}_{t}^{(i)}$. Therefore, Theorem 6.3 immediately gives a bound for Definition 6.8 since
(48) $\|\mathbf{A}^{-1/2}\mathbf{B}-\mathsf{bisq}_{t}\|$ $=\|\mathbf{A}^{-1/2}\mathbf{B}-\mathbf{A}^{-1}\mathsf{bsq}_{t}\|$
(49) $=\|\mathbf{A}^{-1}(\mathbf{A}^{1/2}\mathbf{B}-\mathsf{bsq}_{t})\|$
(50) $\leq\|\mathbf{A}^{-1}\|\|\mathbf{A}^{1/2}\mathbf{B}-\mathsf{bsq}_{t}\|.$
@@ 7 Numerical experiments
We use the block-Lanczos algorithm with full-reorthogonalization in order to implement block-CG and the block-Lanczos square root iterates. Code to replicate the experiments in this paper is available at https://github.com/tchen-research/precond_without_precond. The $s$ matrix-loads required to build the Nyström preconditioner are accounted for in the plots. For Nyström PCG, we set $\theta$ as the smallest eigenvalue of the Nyström approximation. We use the same random Gaussian matrix $\mathbf{\Omega}\in\mathbb{R}^{d\times\ell}$ for Nyström PCG and block-CG. More numerical experiments, including without full-reorthogonalization, are included in Appendices B and A.
@@ 7.1 Convergence
We begin by comparing the convergence of the methods discussed in this paper on several test problems. The results of this experiment are illustrated in Fig. 2, and as expected, our augmented block-CG method outperforms the other methods, often by orders-of-magnitude. We also observe that Nyström PCG benefits from using $s>1$; i.e. from performing more than one passes over $\mathbf{A}$ when building the preconditioner.
As discussed in Section 1.2, throughout this paper we assume that matrix-loads (iterations) are the dominant cost. For reference, we have also show convergence as a function of matrix-vector products in Fig. 8.
@@ 7.1.1 Intro figure
Figure 1 is run on the fastdecay problem ($d=8000$) with $\ell=10$. Nyström preconditioning is run with a grid of parameters. Specifically, we use depths $s\in\{1,5,11\}$ and a range of $\theta$, including the choice $\theta=\lambda_{\ell}(\mathbf{A}\langle\mathbf{\Omega}\rangle)$.
The timings are produced based on a setting where $\mathbf{A}$ cannot be stored in its entirety in fast memory. In particular, we store $\mathbf{A}$ in 8 $1000\times 8000$ chunks and perform (block) matrix-vector products with $\mathbf{A}$ by sequentially loading a single chunk from the disk into random access memory and performing the appropriate part of the products. Similar to [61, §6.1], we observe that the runtime of the algorithm is dominated by the cost of loading these chunks into random access memory.
@@ 7.2 Regularization parameter
In Fig. 3 we plot the error after a fixed number of matrix loads as a function of $\mu$. As expected, this plot indicates that our augmented block-CG outperforms the other methods for each value of $\mu$. However, block-CG reaches its maximum attainable accuracy sooner, stagnating.
Perhaps more importantly, our block-CG method can efficiently compute the solution to Eq. 1 for many values of $\mu$, without the need for additional matrix-loads (see Theorem 3.7). This is in contrast to Nyström PCG which requires a new run for each value of $\mu$. For tasks such as ridge regression, this is of note.
@@ 7.3 Sampling Gaussians
We perform an experiment comparing the block-Lanczos square root iterate to the standard Lanczos square root iterate. In particular, in Fig. 4, we compare the maximum error in approximating $\mathbf{A}^{1/2}\mathbf{b}^{(i)}$, $i=1,\ldots,m$ for the two methods. All runs of the single-vector method are done in parallel so that the number of matrix-loads is independent of $m$. The block method outperforms the single-vector method, and the performance gains are often significant.
@@ 8 Outlook
We have introduced a variant of the block-CG method which outperforms Nyström PCG in certain settings. Our work provides theoretical evidence of the virtues of block Krylov subspace methods for solving a single linear system of equations and sampling Gaussian vectors. We believe future work should study a more practical variant of augmented block-CG, where deflation is used to reduce the block-size as soon as it is recognized that the benefits of a large block size are no longer justified.
It would also be interesting to understand the extent to which similar ideas can be extended to other tasks involving matrix-functions, such as estimating the trace of matrix-functions. Existing algorithms often involve applying matrix-functions to many Gaussian vectors [62, 21, 44], so it seems likely that using block-Krylov methods may be advantageous to using single-vector methods. However, existing theory does not reflect this.
@@ A.1 Convergence
We provide more test problems comparing the convergence of block-CG with Nyström PCG and CG as described in Section 7.1. As observed in Fig. 2, block-CG outperforms the other methods in terms of matrix-loads, and often significantly so.
Note that on the outliers20 problem we now append $\ell=10$ or $\ell=22$ Gaussian vectors (rather than $\ell=20$ as shown in Fig. 2). This problem has 20 large eigenvalues, and the small amount of oversampling significantly improves the convergence of Nyström PCG with $s=1$. We also include a test with the bottom20 problem, which has 20 small eigenvalues.
Figure 5: Relative error $\|\mathbf{A}^{-1}\mathbf{b}-\mathsf{alg}\|_{\mathbf{A}}/\|\mathbf{A}^{-1}\mathbf{b}\|_{\mathbf{A}}$ versus matrix-loads for block-CG (Refer to caption), CG (Refer to caption), and Nyström PCG with $s=1$ (Refer to caption) and $s=3$ (Refer to caption) on several test problems; see also Fig. 2.
@@ A.2 Block size
Theorem 4.3 suggests the accuracy of low-rank approximations to $\mathbf{A}$ built from the information in the block Krylov subspace $\mathcal{K}_{t}(\mathbf{A},\mathbf{\Omega}_{b})$ depends on $bk$, regardless of the individual value of $b$ and $k$. In Fig. 6, we show the error (after a fixed number of matrix-loads) as a function of the block size, and in Fig. 7 we show the condition number of $\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2}$ as a function of the block size. In some cases Nyström PCG with depth $s=3$ behaves similarly to Nyström PCG with depth $s=1$ when the same number of matrix-vector products are used. For instance, on the outliers20 problem, both methods have a significant drop in the error (due to a significant drop in the condition number) when dimension of the Krylov subspace is roughly equal to $20$, the number of outlying eigenvalues.
@@ A.3 Matrix-vector cost
In Fig. 8 we show convergence as a function of matrix-vector products on the same test problems as in Fig. 2. Here our block-CG underperforms the other methods. This does not conflict with any of our theory, which is in terms of matrix-loads. However, it serves as a reminder that our method may not be suitable in all computational settings; see Section 1.2.
Figure 8: Relative error $\|\mathbf{A}^{-1}\mathbf{b}-\mathsf{alg}\|_{\mathbf{A}}/\|\mathbf{A}^{-1}\mathbf{b}\|_{\mathbf{A}}$ versus matrix-products for block-CG (Refer to caption), CG (Refer to caption), and Nyström PCG with $s=1$ (Refer to caption) and $s=3$ (Refer to caption) on the same test problems as Fig. 2.
@@ A.4 Test problems
In Fig. 9 we show the spectrums of the test problems we used.
@@ B.1 Bounds for CG in terms of the deflated condition number
In Theorem 4.7 we show that if $s=O(\log(d))$, then $\kappa(\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2})=O(\lambda_{r+1}/\lambda_{d})$. It is informative to see bounds for CG in terms of $\lambda_{r+1}/\lambda_{d}$ as well as in terms of the effective dimension. Towards this end, we recall a standard bound for CG; see e.g. [29].
@@
Theorem B.1 guarantees that CG converges at a rate of $\sqrt{\lambda_{r+1}/\lambda_{d}}$ after a burn-in period of $r$ iterations. This is reminiscent of the bounds for Nyström PCG presented in this paper, which guarantee convergence at this rate after a preconditioner build time of roughly $O(r)$ matrix-products.
@@ B.2 Bounds in terms of the effective dimension
(51) $d_{\textup{eff}}(\mu):=\operatorname{tr}\bigl{(}\mathbf{A}\mathbf{A}_{\mu}^{-1}\bigr{)}=\sum_{i=1}^{d}\frac{\lambda_{i}}{\lambda_{i}+\mu}.$
@@
When $\kappa(\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2})$ is bounded by a constant, preconditioned CG will converge to a constant accuracy (e.g. $10^{-4}$ or $10^{-16}$) in a number of iterations independent of the condition number or spectral properties of $\mathbf{A}$.
We can convert Theorem B.1 to a bound in terms of the effective dimension by noting that all but the first $O(d_{\textup{eff}})$ eigenvalues of $\mathbf{A}$ are bounded by $\mu$. In particular, [23, Lemma 5.4] implies
@@
Together, Theorems B.1 and 53 give a bound for CG in terms of the effective dimension.
@@
The CG iterate satisfies
@@
Corollary B.3 asserts that after a burn in of $2d_{\textup{eff}}(\mu)$ iterations, CG converges in $O(\log(1/\varepsilon))$ iterations. Thus, the theory in [23] does not guarantee Nyström PCG has any major benefits over CG in terms of matrix-vector products.
@@ B.3 Is Nyström PCG a good idea?
The existence of a bound like Theorem B.1 seems damning; CG automatically satisfies bounds similar to those of of Nyström PCG, without the need to construct and store a preconditioner. However, as we now discuss, the full story is much more subtle, and we believe Nyström PCG is still a viable method in many instances.
First, Nyström PCG is able to parallelize matrix-vector products used to build the preconditioner. Thus, the number of matrix-loads can be considerably less than required by CG. This is reflected in our experiments in Fig. 2.
However, these same plots indicated that in terms of matrix-products, CG significantly outperforms Nyström PCG (the $s\ell$ matrix-products to build the Nyström preconditioner require only one matrix-load). On the other hand, in apparent contradiction to our experiments, the experiments in [23] indicate that Nyström PCG consistently and significantly outperforms CG.
Recall that the experiments in Section 7 are done using full-reorthogonalization (so that exact arithmetic theory is still applicable). On the other hand, the experiments in [23] appear to have been done without reorthogonalzation. It’s well-known that Theorem B.1 does not hold finite precision arithmetic without reorthogonalization for any r>0r>0.666A weaker version does still hold [28]. On the other hand, Corollary 2.3 can be expected to hold to close degree [28, 42]. Since Nyström PCG works by explicitly decreasing the condition number of $\kappa(\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2})$, the convergence of Nyström PCG in finite precision arithmetic is still well described by the exact arithmetic theory for Nyström PCG.
@@
We illustrate this point of comparison explicitly in Fig. 10(a). We run CG and Nyström PCG on the fastdecay problem (for which $\lambda_{1}/\lambda_{d}\gg\lambda_{20}/\lambda_{d}$). As expected, CG with reorthogonalization burns in for roughly 20 iterations, and then begins converging much more rapidly at a rate depending on $\lambda_{21}/\lambda_{d}$, outperforming Nyström PCG with reorthogonalization ($s=1,\ell=22$), which uses 22 matrix-vector products [^7] We need some oversampling to capture the top eigenspace. to build the preconditioner, at which point also converges at a rate depending on $\lambda_{21}/\lambda_{d}$. However, if reorthogonalization is not used, CG converges at a much slower rate depending on $\lambda_{1}/\lambda_{d}$ while Nyström PCG still converges at the faster rate depending on $\lambda_{21}/\lambda_{d}$.

### v3

@@ 5.1 Sampling with the inverse covariance
Computing $\mathbf{A}^{-1/2}\mathbf{b}$ is used to transform vectors with covariance matrix $\mathbf{A}$ to “whitened” coordinate space, where the the covariance is the identity, and has found use in a number of data-science applications [42, 59]. We can define an approximation for applying inverse square roots similar to the block-Lanczos square root iterate (˜5.2).
@@
Note that $\mathsf{bisq}_{t}^{(i)}=\mathbf{A}^{-1}\mathsf{bsq}_{t}^{(i)}$. Therefore, ˜5.3 immediately gives a bound for ˜5.7 since
$\|\mathbf{A}^{-1/2}\mathbf{B}-\mathsf{bisq}_{t}\|$ $=\|\mathbf{A}^{-1/2}\mathbf{B}-\mathbf{A}^{-1}\mathsf{bsq}_{t}\|$
$=\|\mathbf{A}^{-1}(\mathbf{A}^{1/2}\mathbf{B}-\mathsf{bsq}_{t})\|$
$\leq\|\mathbf{A}^{-1}\|\|\mathbf{A}^{1/2}\mathbf{B}-\mathsf{bsq}_{t}\|.$
6 Discussion and comparison with past work
There are may other related approaches to solving the problems in Section˜1. For instance, KSMs are commonly used to solve Tikahonov Regression problems, using the shift-invariance of Krylov subspaces Equation˜5 solve for to whole regularization path efficiently [39, 28, 5]. Similarly, as discussed in Section˜5, KSMs are also for tasks involving matrix functions, including sampling Gaussian vectors [13, 59, 8]. The techniques discussed in Section˜5 are likely applicable other functions as well. The block-CG algorithm is widely used to solve linear systems with multiple right hand sides, and in such settings working over the block Krylov subspace is often advantageous compared to working over the individual Krylov subspaces [54, 47, 24, 4]. Block-CG is also used to solve single linear systems. For instance, so-called enlarged KSMs split the right hand size vector $\mathbf{b}$ into multiple vectors using a domain decomposition approach, and then apply block-CG to this collection of vectors [34, 35, 36]. Augmenting a KSM with random vectors was shown to be beneficial in numerical experiments appearing in concurrent work [74, Appendix B]. While we are unaware of any bounds similar to Theorem˜4.3, bounds similar to Theorem˜4.2 are known. In particular, [54, §4] derives a convergence bound for block-CG terms of $\sqrt{\lambda_{1}/\lambda_{d-m+1}}$; i.e. the condition number of $\mathbf{A}$ with the bottom $m$ eigenvalues removed. [^6] Note that the ordering of the eigenvalues in [54] is reversed from the present paper. At a high level, [54] obtains this bound by showing that the block Krylov subspace contains vectors which can approximately annihilate a subset of the eigenvalues of $\mathbf{A}$. Thus, in principle, other sets of eigenvalues (besides the bottom ones) can be eliminated. The bounds of [54] work for any starting block (but as a result, are more technical, depending on how “good” the starting block is). It would not be surprising such quantities can be bounded when the starting block is Gaussian. KSMs are widely used for both low-rank approximation and problems involving matrix-functions, but their behavior on these problems is fundamentally different. As such, it would be interesting to try to avoid relying on a black-box reduction to low-rank approximation. For example, while low-rank approximation is complicated by the presence of repeated eigenvalues, solving linear systems is not. The approach of [54] is closely related to the approaches used in [68, 51] to prove bounds for low-rank approximation, so it is conceivable that, by taking the approach of [54], but using modern tools from randomized numerical linear algebra, one may be able to directly derive better bounds for randomized versions of block-CG. More broadly, there are many randomized algorithms for linear systems and regression. For instance, a recent line of work shows that certain stochastic iterative methods enjoy an implicit preconditioning-like effect [15, 44, 18, 19, 17], with convergence independent of the top eigenvalues. These algorithms have runtime guarantees better than KSMs in some settings [17]. In addition, for tall least squares problems, the sketch-and-precondition framework for building a preconditioner for KSMs is extremely effective [61, 46], resulting in algorithms with optimal theoretical complexity [14, 12]. A number of recent works in randomized numerical linear algebra [11, 68, 50, 9] have made use of various nestedness properties of block-KSMs similar in flavor to Theorem˜3.2 in order to obtain more efficient algorithms.
@@ 7 Numerical experiments
[…]
@@ 7.1 Convergence
We begin by comparing the convergence of the methods discussed in this paper on several test problems. The results of this experiment are illustrated in ˜2, and as expected, our augmented block-CG method outperforms the other methods, often by orders-of-magnitude. We also observe that Nyström PCG benefits from using $s>1$; i.e. from performing more than one passes over $\mathbf{A}$ when building the preconditioner. As discussed in Subsection˜1.2, throughout this paper we assume that matrix-loads (iterations) are the dominant cost. For reference, we have also show convergence as a function of matrix-vector products in ˜8.
@@ 7.1.1 Intro figure
Figure˜1 is run on the fastdecay problem ($d=8000$) with $\ell=10$. Nyström preconditioning is run with a grid of parameters. Specifically, we use depths $s\in\{1,5,11\}$ and a range of $\theta$, including the choice $\theta=\lambda_{\ell}(\mathbf{A}\langle\mathbf{\Omega}\rangle)$. The timings are produced based on a setting where $\mathbf{A}$ cannot be stored in its entirety in fast memory. In particular, we store $\mathbf{A}$ in 8 separate $1000\times 8000$ chunks and perform (block) matrix-vector products with $\mathbf{A}$ by sequentially loading a single chunk from the disk into random access memory and performing the appropriate part of the products. Similar to [68, §6.1], we observe that the runtime of the algorithm is dominated by the cost of loading these chunks into random access memory.
@@ 7.2 Regularization parameter
In ˜3 we plot the error after a fixed number of matrix loads as a function of $\mu$. As expected, this plot indicates that our augmented block-CG outperforms the other methods for each value of $\mu$. Perhaps more importantly, our block-CG method can efficiently compute the solution to Equation˜1 for many values of $\mu$, without the need for additional matrix-loads (see Theorem˜3.7). This is in contrast to Nyström PCG which requires a new run for each value of $\mu$. For tasks such as ridge regression, this is of note. We note the maximum attainable accuracy of block-CG (an effect of performing the numerical experiments in finite precision arithmetic which prevents convergence to the true solution) seems to be higher than that of other methods, causing stagnation.
@@ 7.3 Sampling Gaussians
We perform an experiment comparing the block-Lanczos square root iterate to the standard Lanczos square root iterate. In particular, in ˜4, we compare the maximum error in approximating $\mathbf{A}^{1/2}\mathbf{b}^{(i)}$, $i=1,\ldots,m$ for the two methods. All runs of the single-vector method are done in parallel so that the number of matrix-loads is independent of $m$. The block method outperforms the single-vector method, and the performance gains are often significant.
@@ 8 Outlook
We have introduced a variant of the block-CG method which outperforms Nyström PCG in certain settings. Our work provides theoretical evidence of the virtues of block Krylov subspace methods for solving a single linear system of equations and sampling Gaussian vectors. We believe future work should study a more practical variant of augmented block-CG, where deflation is used to reduce the block-size as soon as it is recognized that the benefits of a large block size are no longer justified. It would also be interesting to understand the extent to which similar ideas can be extended to other tasks involving matrix-functions, such as estimating the trace of matrix-functions. Existing algorithms often involve applying matrix-functions to many Gaussian vectors [69, 23, 50], so it seems likely that using block-Krylov methods may be advantageous to using single-vector methods. However, existing theory does not reflect this.
@@ A.1 Convergence
We provide more test problems comparing the convergence of block-CG with Nyström PCG and CG as described in Subsection˜7.1. As observed in ˜2, block-CG outperforms the other methods in terms of matrix-loads, and often significantly so. Note that on the outliers20 problem we now append $\ell=10$ or $\ell=22$ Gaussian vectors (rather than $\ell=20$ as shown in ˜2). This problem has 20 large eigenvalues, and the small amount of oversampling significantly improves the convergence of Nyström PCG with $s=1$. We also include a test with the bottom20 problem, which has 20 small eigenvalues.
Figure 5: Relative error $\|\mathbf{A}^{-1}\mathbf{b}-\mathsf{alg}\|_{\mathbf{A}}/\|\mathbf{A}^{-1}\mathbf{b}\|_{\mathbf{A}}$ versus matrix-loads for block-CG (Refer to caption), CG (Refer to caption), and Nyström PCG with $s=1$ (Refer to caption) and $s=3$ (Refer to caption) on several test problems; see also ˜2.
@@ A.2 Block size
Theorem˜4.3 suggests the accuracy of low-rank approximations to $\mathbf{A}$ built from the information in the block Krylov subspace $\mathcal{K}_{t}(\mathbf{A},\mathbf{\Omega})$, where $\mathbf{\Omega}$ is a random Gaussian matrix, depends on $bk$, regardless of the individual value of $b$ and $k$. In ˜6, we show the error (after a fixed number of matrix-loads) as a function of the block size, and in ˜7 we show the condition number of $\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2}$ as a function of the block size. In some cases Nyström PCG with depth $s=3$ behaves similarly to Nyström PCG with depth $s=1$ when the same number of matrix-vector products are used. For instance, on the outliers20 problem, both methods have a significant drop in the error (due to a significant drop in the condition number) when dimension of the Krylov subspace is roughly equal to $20$, the number of outlying eigenvalues.
@@ A.3 Matrix-vector cost
In ˜8 we show convergence as a function of matrix-vector products on the same test problems as in ˜2. Here our block-CG underperforms the other methods. This does not conflict with any of our theory, which is in terms of matrix-loads. However, it serves as a reminder that our method may not be suitable in all computational settings; see Subsection˜1.2.
Figure 8: Relative error $\|\mathbf{A}^{-1}\mathbf{b}-\mathsf{alg}\|_{\mathbf{A}}/\|\mathbf{A}^{-1}\mathbf{b}\|_{\mathbf{A}}$ versus matrix-vector products for block-CG (Refer to caption), CG (Refer to caption), and Nyström PCG with $s=1$ (Refer to caption) and $s=3$ (Refer to caption) on the same test problems as ˜2.
@@ A.4 Test problems
In ˜9 we show the spectrums of the test problems we used.
@@ B.1 Bounds for CG in terms of the deflated condition number
In Theorem˜4.7 we show that if $s=O(\log(d))$, then $\kappa(\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2})=O(\lambda_{r+1}/\lambda_{d})$. It is informative to compare such a bound against similar bounds for CG. Towards this end, we recall a standard bound for CG; see e.g. [33].
@@
˜B.1 guarantees that CG converges at a rate of $\sqrt{\lambda_{r+1}/\lambda_{d}}$ after a burn-in period of $r$ iterations. This is reminiscent of the bounds for Nyström PCG presented in this paper, which guarantee convergence at this rate after a preconditioner build time of roughly $O(r)$ matrix-products.
@@ B.2 Bounds in terms of the effective dimension
$d_{\textup{eff}}(\mu):=\operatorname{tr}\bigl(\mathbf{A}\mathbf{A}_{\mu}^{-1}\bigr)=\sum_{i=1}^{d}\frac{\lambda_{i}}{\lambda_{i}+\mu}.$
@@
When $\kappa(\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2})$ is bounded by a constant, preconditioned CG will converge to a constant accuracy (e.g. $10^{-4}$ or $10^{-16}$) in a number of iterations independent of the condition number or spectral properties of $\mathbf{A}$. We can convert ˜B.1 to a bound in terms of the effective dimension by noting that all but the first $O(d_{\textup{eff}})$ eigenvalues of $\mathbf{A}$ are bounded by $\mu$. In particular, [25, Lemma 5.4] implies that for any $r$ and $\mu\geq 0$,
@@
[…]
@@
corollary]thm:CG_deff The CG iterate satisfies
@@
LABEL:thm:CG_deff asserts that after a burn in of $2d_{\textup{eff}}(\mu)$ iterations, CG converges in $O(\log(1/\varepsilon))$ iterations. Thus, the theory in [25] does not guarantee Nyström PCG has any major benefits over CG in terms of matrix-vector products.
@@ B.3 Is Nyström PCG a good idea?
The existence of a bound like ˜B.1 seems damning; CG automatically satisfies bounds similar to those of of Nyström PCG, without the need to construct and store a preconditioner. However, as we now discuss, the full story is much more subtle, and we believe Nyström PCG is still a viable method in many instances. First, Nyström PCG is able to parallelize matrix-vector products used to build the preconditioner. Thus, the number of matrix-loads can be considerably less than required by CG. This is reflected in our experiments in ˜2. However, these same plots indicated that in terms of matrix-products, CG significantly outperforms Nyström PCG (the $s\ell$ matrix-products to build the Nyström preconditioner require only one matrix-load). On the other hand, in apparent contradiction to our experiments, the experiments in [25] indicate that Nyström PCG consistently and significantly outperforms CG. Recall that the experiments in Section˜7 are done using full-reorthogonalization (so that exact arithmetic theory is still applicable). On the other hand, the experiments in [25] appear to have been done without reorthogonalzation. It’s well-known that ˜B.1 does not hold finite precision arithmetic without reorthogonalization for any r>0r>0.777A weaker version does still hold [32]. On the other hand, Corollary˜2.3 can be expected to hold to close degree [32, 48]. Since Nyström PCG works by explicitly decreasing the condition number of $\kappa(\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2})$, the convergence of Nyström PCG in finite precision arithmetic is still well described by the exact arithmetic theory for Nyström PCG.
@@
We illustrate this point of comparison explicitly in ˜10(a). We run CG and Nyström PCG on the fastdecay problem (for which $\lambda_{1}/\lambda_{d}\gg\lambda_{20}/\lambda_{d}$). As expected, CG with reorthogonalization burns in for roughly 20 iterations, and then begins converging much more rapidly at a rate depending on $\lambda_{21}/\lambda_{d}$, outperforming Nyström PCG with reorthogonalization ($s=1,\ell=22$), which uses 22 matrix-vector products [^8] We need some oversampling to capture the top eigenspace. to build the preconditioner, at which point also converges at a rate depending on $\lambda_{21}/\lambda_{d}$. However, if reorthogonalization is not used, CG converges at a much slower rate depending on $\lambda_{1}/\lambda_{d}$ while Nyström PCG still converges at the faster rate depending on $\lambda_{21}/\lambda_{d}$.